# Partial Regularity for the three-dimensional Stochastic Ericksen–Leslie equations

### v1

@@
Abstract.
In this article, we investigate the existence of martingale suitable weak solutions to 3D stochastic Ericksen–Leslie equations driven by an external random force. Despite the inherent white noise, we prove that paths of martingale suitable weak solutions exhibit singular points of one-dimensional parabolic Hausdorff measure zero. Our approach is mainly based on two ideas: a local energy-preserving approximation which serves to establish the existence of suitable weak solutions, and a blow-up argument employed to achieve the partial regularity.
1. Introduction
@@
Liquid crystal, an anisotropic phase situated between the isotropic liquid phase and solid phase, has long intrigued scientists from various disciplines. Among many types of liquid crystal states, the nematic phase is particularly notable for its characteristic that in this kind of material, the rod-shaped molecules display the orientational order (tending to point in the same direction) without positional order. The hydrodynamics of nematic liquid crystals has been a subject of extensive research, see for instance [25, 38, 49, 50]. A fundamental mathematical model for describing hydrodynamics in nematic liquid crystal flows is provided by the Ericksen–Leslie equations [16, 17, 33].
@@
The Ericksen–Leslie equations combine momentum balance, typically captured by the Navier–Stokes equation [21, 45, 46], with the evolution of molecular orientation, characterized by the heat flow of harmonic maps [36]. In this article, we focus on the simplified version of the full Ericksen–Leslie equations, which was initially studied by Lin and Liu [34, 35], driven by a random force:
@@
(1.1) $\left\{\begin{array}[]{l}\,\mathrm{d}{{\bf{u}}}-\Delta{{\bf{u}}}\,\mathrm{d}t+{{\bf{u}}}\cdot\nabla{{\bf{u}}}\,\mathrm{d}t+\nabla P\,\mathrm{d}t=-\mathop{\rm div}\nolimits(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}})\,\mathrm{d}t+\,\mathrm{d}W,\\ \mathop{\rm div}\nolimits{{\bf{u}}}=0,\\ \,\mathrm{d}{{\bf{d}}}-\Delta{{\bf{d}}}\,\mathrm{d}t+{{\bf{u}}}\cdot\nabla{{\bf{d}}}\,\mathrm{d}t=-f_{\epsilon}({{\bf{d}}})\,\mathrm{d}t.\end{array}\right.$
@@
Here the unknown triple $({{\bf{u}}},P,{{\bf{d}}})$ represents the fluid velocity field, the pressure function, and the director field which characterizes the mean orientation of molecules, and $f_{\epsilon}({{\bf{d}}})=\nabla_{{\bf{d}}}F_{\epsilon}({{\bf{d}}})$, where
$F_{\epsilon}({{\bf{d}}})=\frac{1}{\epsilon^{2}}(|{{\bf{d}}}|^{2}-1)^{2}.$
It is evident that (1.1)${}_{1,2}$ is a forced Stokes system with the so-called Ericksen stress tensor $(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}})_{ij}=\partial_{x_{i}}{{\bf{d}}}\cdot\partial_{x_{j}}{{\bf{d}}}$ and a random force $W$ which will be specified later, while (1.1)${}_{3}$ is a transported semi-linear heat flow associated to the Ginzburg–Landau potential function $F_{\epsilon}({{\bf{d}}})$. The ultimate goal is to study the singular limit as $\epsilon\to 0$ in (1.1)${}_{3}$, which leads to the following transported heat flow of harmonic maps for the director field ${\bf{d}}$:
@@
It is well-established that the critical dimension for heat flow of harmonic maps is two [8, 9]. Consequently, in 2D, it has been confirmed that solutions to the deterministic (limited) Ericksen–Leslie equation are partially regular [37], and singularities could occur [26, 30]. Moreover, the singular limit as $\epsilon\to 0$ has been justified through concentration-cancellation compactness [29], as also discussed in [14]. However, in 3D, the well-posedness for the deterministic (limited) Ericksen–Leslie equation remains an open problem with only a few partial results available, such as those presented in [39, 47].
@@
Despite vast success in studies of the deterministic Ericksen–Leslie equations, it is worth noting that real-world scenarios introduce complexities that a deterministic framework could possibly overlook, including environmental disturbances [10], measurement uncertainties, and inherent thermal fluctuations [3, 31]. To address these intricacies, a typical way is to introduce a stochastic component into the Ericksen–Leslie equations, which leads to the formulation of the stochastic Ericksen–Leslie (SEL) equations. A probable consequence of the stochastic force in the nematic liquid crystal flows is its potential impact on the emergence of singularities, which can manifest as either turbulence in fluid flow or defects in molecular alignments. These phenomena carry substantial implications, impacting our understanding of theory and the practical applications of liquid crystal materials.
To this end, in this article, we study the SEL equations on the three-dimensional tori $\mathbb{T}^{3}$ driven by additive white noise. Despite a recent growing interest in the SEL equations, to our best knowledge, there have been no reported results regarding the regularity of the SEL equations in three dimensions. This research gap persists even when $\epsilon$ in (1) is held fixed rather than tending towards zero, as mentioned in [4, 5]. It’s worth mentioning that well-posedness results concerning the (limited) SEL equations in two dimensions can be found in [12] (with additive noise), and in [48] (with multiplicative noise); for further details, one may refer to [6, 16, 44]. Here our main objectives are two-fold. Firstly, we introduce the concept and establish the existence of martingale suitable weak solutions to (1.1). Since $\epsilon$ plays no role in our analysis, from now on we fix $\epsilon=1$ and omit it from $F_{\epsilon}$ and $f_{\epsilon}$. It is worth noting that the concept of martingale solutions is equivalent to that of statistical solutions, a topic extensively studied in statistical fluid dynamics (see for instance [22, 23, 24]). Secondly, we demonstrate that, with certain assumptions on the white noise, the set of singular points, where solutions became unbounded, cannot have a positive one-dimensional parabolic Hausdorff measure.
The article is organized as follows. In Section 2, we specify notations and stochastic framework to set up the SEL equations. Then in Section 3 we introduce a notion of suitable weak solutions, drawing inspiration from [20, 42] in the context of the stochastic Navier–Stokes equation. This approach involves splitting solutions to the SEL equations into two parts. The first part is governed by linear stochastic Stokes equations, and the second part is subject to the remaining nonlinear coupled system. Then in Section 4, we prove the existence of martingale suitable weak solutions. Finally in Section 5 we establish the partial regularity of martingale suitable weak solutions.
2. Mathematical settings and notations
2.1. Functional spaces
@@
We introduce two fundamental spaces that take into account the incompressibility condition (1.1)${}_{2}$ (c.f. [46]):
@@
The operator $A:D(A)\subset H\to H$ is defined as $Au=-P_{L}\Delta{\bf{u}}$, where $P_{L}$ is the Helmholtz–Leray projection from $L^{2}(\mathbb{T}^{3};\mathbb{R}^{3})$ onto $H$ and $D(A)=H^{2}(\mathbb{T}^{3};\mathbb{R}^{3})\cap V$. It is well-known that $A^{-1}$ is well-defined from $H$ to $D(A)$. It is also self-adjoint. By the spectral theory of compact self-adjoint operators in a Hilbert space, there exist a sequence of eigenvalues $0<\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{m}\leq\cdots$, $\lambda_{m}\to\infty$ as $m\to\infty$ and an orthonormal basis $\{e_{i}\}_{i=1}^{\infty}$ such that $Ae_{i}=\lambda_{i}e_{i}$. The fractional power $A^{\alpha}$ of $A$, $\alpha\geq 0$, are simply defined by
@@
Also we note that $D(A^{\alpha})\subset H^{2\alpha}(\mathbb{T}^{3};\mathbb{R}^{3})$. For the definition of $H^{2\alpha}(\mathbb{T}^{3};\mathbb{R}^{3})$, we refer to [41, Definition 1.10].
2.2. Stochastic framework
@@
(2.1) $\mathfrak{B}:=\left(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P},\left(W(t)\right)_{t\geq 0}\right),$
@@
(2.2) $\left(W(t)\right)_{t\geq 0}$ takes value in $D(A^{\delta})$ for some $\delta>0$ with a covariance operator $\mathcal{L}$.
Here $\mathcal{L}\in\mathcal{B}(D(A^{\delta}))$ and is non-negative and of trace class. Let $\{b_{i}\}_{i\geq 1}$ be the orthonormal basis of $D(A^{\delta})$ consisting of eigenvectors of $\mathcal{L}$ subject to non-negative eigenvalues $\{\gamma_{i}\}_{i\geq 1}$. Then we have the following expansion (c.f. [40, Theorem 4.20])
@@
(2.3) $W(t)\in C^{\beta}\left([0,T];D(A^{\delta})\right)\quad\mathbb{P}-a.s..$
3. Martingale suitable weak solution
@@
We will assume throughout the paper the following assumptions for the initial data
(3.1) $({{\bf{u}}}_{0},{{\bf{d}}}_{0})\in H\times H^{1}(\mathbb{T}^{3};\mathbb{R}^{3}),\quad|{{\bf{d}}}_{0}(x)|=1\text{ for all }x\in\mathbb{T}^{3}.$
@@
As a preparation to introduce notions of deterministic and stochastic suitable weak martingale solutions, we will first split the solutions ${{\bf{u}}}$ and $P$ properly. We introduce new variables ${{\bf{u}}}={{\bf{z}}}+{{\bf{v}}}$, $P={Q}+{\pi}$ where $({{\bf{z}}},{Q})$ solves the linear stochastic Stokes system as follows
@@
(3.2) $\left\{\begin{array}[]{l}{\,\mathrm{d}}{{\bf{z}}}-\Delta{{\bf{z}}}{\,\mathrm{d}}t+\nabla{Q}{\,\mathrm{d}}t={\,\mathrm{d}}W,\\ \mathop{\rm div}\nolimits{{\bf{z}}}=0,\\ {{\bf{z}}}(0)=0.\end{array}\right.$
While the triple $({{\bf{v}}},{\pi},{{\bf{d}}})$ solves the modified Ericksen–Leslie equations as follows
@@
(3.3) $\left\{\begin{array}[]{l}\partial_{t}{{\bf{v}}}-\Delta{{\bf{v}}}+({{\bf{z}}}+{{\bf{v}}})\cdot\nabla({{\bf{z}}}+{{\bf{v}}})+\nabla{\pi}=-\mathop{\rm div}\nolimits\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}-F({{\bf{d}}}){\rm I}_{3}\right),\\ \mathop{\rm div}\nolimits{{\bf{v}}}=0,\\ \partial_{t}{{\bf{d}}}-\Delta{{\bf{d}}}+({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{d}}}=-f({{\bf{d}}}),\\ {{\bf{v}}}(0)={{\bf{u}}}_{0},{{\bf{d}}}(0)={{\bf{d}}}_{0}.\end{array}\right.$
Remark 3.1.
@@
Due to the incompressibility condition (3.3)${}_{2}$, the right-hand side of (3.3)${}_{1}$ is equivalent to (${}_{1}$, since for the extra term we have
$\displaystyle\mathop{\rm div}\nolimits\left(\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}+F({{\bf{d}}}){\rm I}_{3}\right)=\nabla\left(\frac{1}{2}|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right),$
@@
which then can be treated as part of the pressure gradient. On the other hand, direct computations yield
@@
(3.4) $-\mathop{\rm div}\nolimits\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}-F({{\bf{d}}}){\rm I}_{3}\right)=-\nabla{\bf{d}}\cdot(\Delta{\bf{d}}-f({\bf{d}})).$
Let ${{\bf{z}}}(t)$ be the stochastic process that solves (3.2). In the mild form we have
By the Doob inequality and the Burkholder–Davis–Gundy inequality, for $p>1$, $\delta>0$, and $\varepsilon>0$ we can estimate
(3.5) $\mathbb{E}\left[\sup_{0\leq t\leq T}\|{{\bf{z}}}(t)\|_{D(A^{\delta+\frac{1}{2}-\varepsilon})}^{p}\right]$ $=\mathbb{E}\left[\sup_{0\leq t\leq T}\|A^{\delta+\frac{1}{2}-\varepsilon}{{\bf{z}}}(t)\|_{L^{2}(\mathbb{T}^{3})}^{p}\right]$
$\leq C(p)\mathbb{E}\left(\sum_{i=1}^{\infty}\gamma_{i}\int_{0}^{T}\|A^{\frac{1}{2}-\varepsilon}e^{-(T-s)A}A^{\delta}b_{i}\|_{L^{2}(\mathbb{T}^{3})}^{2}\,\mathrm{d}s\right)^{\frac{p}{2}}$
$\leq C(p)\mathbb{E}\left(\sum_{i=1}^{\infty}\gamma_{i}\int_{0}^{T}\|A^{\frac{1}{2}-\varepsilon}e^{-(T-s)A}\|^{2}_{\mathcal{B}(L^{2}(\mathbb{T}^{3}))}\|A^{\delta}b_{i}\|_{L^{2}(\mathbb{T}^{3})}^{2}\,\mathrm{d}s\right)^{\frac{p}{2}}$
[…]
where $C(p)$ is a constant that only depends on $p$. This implies that for a.s. $\omega$, ${{\bf{z}}}(\omega)\in L^{\infty}(0,T;D(A^{\delta+\frac{1}{2}-\varepsilon}))$. In particular, if we set $\delta=\frac{1}{4}+\beta$, $\beta>\varepsilon$, then by the Sobolev embedding
$D(A^{\frac{3}{4}+\beta-\varepsilon})\subset H^{\frac{3}{2}+2\beta-2\varepsilon}(\mathbb{T}^{3})\subset L^{\infty}(\mathbb{T}^{3}),$
based on which we have the following lemma which ensures the boundedness of ${{\bf{z}}}(\omega)$:
Lemma 3.2.
[20, Lemma 2.2] Fix $\omega\in\Omega$ with $W(\omega)\in C^{\frac{1}{2}-\varepsilon}([0,\infty);D(A^{\frac{1}{4}+\beta}))$ for some $\beta>\varepsilon>0$, then
(3.6) ${{\bf{z}}}(\omega)\in L_{\rm loc}^{\infty}(\mathbb{T}^{3}\times[0,\infty)).$
Definition 3.1.

### v2

@@
Abstract
In this article, we investigate the global existence of martingale suitable weak solutions to stochastic Ericksen–Leslie equations with additive noise in a 3D torus. The notion of suitable weak solutions has been introduced to address possible emergence of finite-time singularities, which remains a notably challenging question in the field of fluid dynamics. Weak solutions offer an approach to account for these potential singularities. A restricted class of weak solutions that exhibit a higher level of regularity which are therefore more likely to be physically meaningful, is naturally called for. Consequently, suitable weak solutions, i.e., weak solutions that satisfy a local energy inequality, become a focus of research, including investigations about how regular these solutions can be. In this article, we prove that, despite the presence of white noise, the paths of martingale suitable weak solutions of 3D stochastic Ericksen–Leslie equations exhibit singular points of one-dimensional parabolic Hausdorff measure zero. To establish this result, we have utilized two techniques, which can potentially be generalized to handle other stochastically forced complex fluid dynamics equations with a similar structure. Firstly, a local energy-preserving approximation is constructed which markedly facilitates the proof of the global existence of martingale suitable weak solutions; secondly, to demonstrate partial regularity of these solutions, a blow-up argument is formulated, which efficiently yields the desired key estimate.
keywords:
Stochastic PDEs, Partial regularity, Complex fluid dynamics, Hydrodynamics of liquid crystal
pacs:
MSC Classification]60H15, 35R60, 60H30, 76D06, 76D03
1 Introduction
@@
Liquid crystal, an anisotropic phase situated between the isotropic liquid phase and solid phase, has drawn considerable attention from researchers from various disciplines. Among all types of liquid crystal states, the nematic phase is particularly interesting for its characteristic that rod-shaped molecules display the orientational order (tending to point in the same direction) without positional order. The Ericksen–Leslie equations (see for instance [22, 23, 61]) serve as a fundamental mathematical model for describing the hydrodynamics of nematic liquid crystal flows, a subject of extensive research by itself [68, 91, 92]. To characterize orientations of nematic liquid crystal molecules, ‘order parameters’ are commonly adopted; the so-called ‘director theory’, which covers the Ericksen–Leslie theory [62] as a special case, employs unit vector field to represent alignment of the molecules.
@@
The Ericksen–Leslie equation system typically combines momentum balance with the evolution of molecular orientation, captured by the Navier–Stokes equations [27, 86, 87] and the heat flow of harmonic maps [66], respectively. In this article we focus on the simplified version of the full Ericksen–Leslie equation system, which was initially studied in a series of works including [64, 65]:
@@
$\left\{\begin{array}[]{l}\,\mathrm{d}{{\bf{u}}}^{\epsilon}-\Delta{{\bf{u}}}^{\epsilon}\,\mathrm{d}t+{{\bf{u}}}^{\epsilon}\cdot\nabla{{\bf{u}}}^{\epsilon}\,\mathrm{d}t+\nabla P^{\epsilon}\,\mathrm{d}t=-\mathop{\rm div}\nolimits(\nabla{{\bf{d}}}^{\epsilon}\odot\nabla{{\bf{d}}}^{\epsilon})\,\mathrm{d}t,\\ \mathop{\rm div}\nolimits{{\bf{u}}}^{\epsilon}=0,\\ \,\mathrm{d}{{\bf{d}}}^{\epsilon}-\Delta{{\bf{d}}}^{\epsilon}\,\mathrm{d}t+{{\bf{u}}}^{\epsilon}\cdot\nabla{{\bf{d}}}^{\epsilon}\,\mathrm{d}t=-f_{\epsilon}({{\bf{d}}}^{\epsilon})\,\mathrm{d}t.\end{array}\right$. (1)
@@
Here with every fixed $\epsilon>0$, the unknown triple $({{\bf{u}}}^{\epsilon},P^{\epsilon},{{\bf{d}}}^{\epsilon})$ represents the fluid velocity field, the pressure function, and the director field which characterizes the mean orientation of molecules, respectively. Moreover
$f_{\epsilon}({{\bf{d}}}^{\epsilon})=\nabla_{{\bf{d}}}F_{\epsilon}({{\bf{d}}}^{\epsilon}),$ (2)
where
$F_{\epsilon}({{\bf{d}}}^{\epsilon})=\frac{1}{\epsilon^{2}}(|{{\bf{d}}}^{\epsilon}|^{2}-1)^{2}.$ (3)
It is evident that (1)1,2 is a forced Stokes system with the so-called Ericksen stress tensor
$(\nabla{{\bf{d}}}^{\epsilon}\odot\nabla{{\bf{d}}}^{\epsilon})_{ij}=\partial_{x_{i}}{{\bf{d}}^{\epsilon}}\cdot\partial_{x_{j}}{{\bf{d}}}^{\epsilon},$
while (1)3 is a transported semi-linear heat flow associated to the Ginzburg–Landau potential function $F_{\epsilon}({{\bf{d}}}^{\epsilon})$.
Subsequently we will only consider (1) with a fixed $\epsilon$; without loss of generality we will fix $\epsilon=1$ and omit the corresponding subscripts and superscripts. Indeed exploration of the singular limit as $\epsilon\to 0$ in (1)3 is an ultimate goal, which leads to the transported heat flow of harmonic maps for the director field ${\bf{d}}$ as follows:
@@
It is well-established that the critical spatial dimension for heat flow of harmonic maps is two [12, 13]. Consequently, in space dimension two, it has been confirmed that solutions to the (limited) Ericksen–Leslie equation are partially regular [67], and singularities could occur [46, 55]. Moreover, in space dimension two, the singular limit as $\epsilon\to 0$ has been justified through concentration-cancellation compactness [20, 54]. However, in space dimension three, the global well-posedness for the (limited) Ericksen–Leslie equation remains a prominent challenge, with only a handful of partial results available, such as those presented in [36, 37, 38, 39, 40, 69, 89].
@@
While the study of deterministic Ericksen–Leslie equation system has attracted considerable attention, it is worth noting that the deterministic framework could possibly overlook complexities introduced by real-world scenarios, including environmental disturbances [14], measurement uncertainties, and inherent thermal fluctuations [6, 57]. To address these intricacies, a typical way is to introduce a stochastic component into the system. Stochastic forcing in the nematic liquid crystal flows may potentially lead to emergence of singularities, which can cause turbulence in fluid flow [10, 31, 44, 85] and defects in molecular alignments [18, 56, 88]. These phenomena bear substantial implications, impacting our understanding of both theory and the practical applications of liquid crystal materials.
To this end, in this article, we study the stochastically forced Ericksen–Leslie equations (SEL) on the three-dimensional tori $\mathbb{T}^{3}$ driven by an additive white noise $W$ as follows:
$\left\{\begin{array}[]{l}\,\mathrm{d}{{\bf{u}}}-\Delta{{\bf{u}}}\,\mathrm{d}t+{{\bf{u}}}\cdot\nabla{{\bf{u}}}\,\mathrm{d}t+\nabla P\,\mathrm{d}t=-\mathop{\rm div}\nolimits(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}})\,\mathrm{d}t+\,\mathrm{d}W,\\ \mathop{\rm div}\nolimits{{\bf{u}}}=0,\\ \,\mathrm{d}{{\bf{d}}}-\Delta{{\bf{d}}}\,\mathrm{d}t+{{\bf{u}}}\cdot\nabla{{\bf{d}}}\,\mathrm{d}t=-f({{\bf{d}}})\,\mathrm{d}t.\end{array}\right$. (4)
Here we note again that $f({{\bf{d}}})$ ccorresponds to $f_{\epsilon}({{\bf{d}}}^{\epsilon})$ when $\epsilon$ is fixed as 1 (see also Equations (2) and (3)). Our main objectives here are two-fold. Firstly, we introduce the concept of martingale suitable weak solutions to (4) and establish their global existence. Secondly, we demonstrate that under certain relatively reasonable assumptions on $W$ (see Section 2.2), the set of singular points, where solutions become unbounded, cannot have a positive one-dimensional parabolic Hausdorff measure. We remark that the concept of martingale solutions that we formulate here is equivalent to that of statistical solutions, a topic extensively studied in fluid dynamics (see for instance [28, 29, 30]).
It is worthwhile to discuss the physical relevance of our choice of modeling the stochastic nematic liquid crystals dynamics by introducing noise into the fluid velocity, as in (4). For example, dynamic scattering, a well-documented and extensively studied physical phenomenon observed in nematics (see, e.g., [34, 35, 52, 72, 43, 48]), typically arises in experiments where the cellular flow of the liquid intensifies to the point of becoming turbulent. Here the liquid velocity plays a central role in the system dynamics, heavily affecting the alignment of the optical director (see e.g. [18]). As a result, it is reasonable to treat liquid flow as the main driver of the system dynamics and assume velocity fluctuations as the dominant source of randomness. Moreover, there are other electrohydrodynamic instabilities identified in previous studies (see e.g. [3, 76, 70, 71, 73, 50]) that also support the class of SEL models in which liquid flow plays a central role. A classical example is the Williams domain instability. Visually resembling the Rayleigh-Bénard instability in isotropic fluids, it can produce a variety of flow geometries depending on the angle between the wave vector and the director field, including normal rolls, oblique rolls, and more complex structures [41, 42, 77]. Additionally, some recent works on stochastic nematic liquid crystal systems also employ this SEL modeling framework where noise is introduced in the velocity component; see, for example, [33, 75].
The question of whether fluid dynamics equations such as 3D Navier-Stokes equations develop finite-time singularities has been a well-known open problem for a long time; to take into account these singularities, it becomes necessary to consider weak solutions, which live in a larger function space than that of continuous or differentiable functions. However, issues may arise with weak solutions, such as non-uniqueness and non-physical properties (see e.g. [2, 10]), due to insufficient regularity. With this motivation, the concept of suitable weak solutions which satisfy a local energy inequality is introduced (for the first time in [60]). Subsequently, a natural question arises: how regular can these solutions be? Various results have been found for the Navier–Stokes equation, as in [80, 81], where suitable weak solutions are shown to be smooth except for a singular set of parabolic Hausdorff dimension $5/3$. Later on in [11], this result was improved; the dimension of this singular set is demonstrated to be no greater than 1. Since then, numerous works have been carried out to explore partial regularity in the context of general fluid dynamics equations, see e.g. [15, 32, 49]. Notably, in [65], a partial regularity result for the deterministic Ericksen–Leslie equations (4) in space-dimension three was achieved.
In contrast, there are relatively limited findings regarding the partial regularity for stochastic PDEs. To the best of our knowledge, such investigations have primarily focused on the stochastic Navier-Stokes equations; notably, in [26, 79], it was established that in 3D the same partial regulairy result as in [11] holds for martingale suitable weak solutions for almost surely every random path. Building upon the insights of [26, 79], here in this article we establish the same type of partial regularity result for SEL in space dimension three. As far as we know, there has been few reported results about the regularity of the SEL equations in space dimension three, even with $\epsilon$ being fixed ([7, 8]), despite a growing interest in this topic recently. In space dimension two, the well-posedness results concerning the SEL equations have been founded in [17] (with additive noise) and [90] (with multiplicative noise), and for further details see e.g. [9, 84].
The article is organized as follows. In Section 2, we specify notations and stochastic framework needed to set up the SEL system. Then in Section 3 we introduce the notion of suitable weak solutions, drawing inspirations from a split-up scheme which is introduced in [26, 79] in the context of the stochastic Navier–Stokes equation. Specifically, this approach involves splitting solutions to the SEL equations into two components. The first part is governed by linear stochastic Stokes equations, and the second part is subject to the remaining nonlinear coupled system. For the first component of the solution, we directly obtain a desired bound, whenever a the noise term is bounded in a suitable fractional Sobolev space almost surely (Section 3.2). Subsequently over each random path that satisfies this assumption, we define deterministic suitable weak solutions which satisfy two local energy inequalities (Section 3.2). Thus we come up with the notion of a martingale suitable weak solution as a martingale solution each trajectory of which is almost surely a deterministic suitable weak solution. Next in Section 3.3, the main result of the article is stated, namely, the global existence and partial regularity of martingale suitable weak solutions. These two claims are proved in Sections 4 and 5, respectively. For the proof of the global existence (Section 4), it is worth noting that the main challenge is to construct an approximation such that the local energy inequality is preserved after passage to the limit. Here we employ a specific type of approximation by regularization, which was introduced in [60] for the Navier–Stokes equations. In [11], a similar approximation is adopted through a ’retarded mollification’ to establish the existence of suitable weak solutions for the Navier–Stokes equations.
2 Mathematical settings and notations
2.1 Functional spaces
@@
We introduce two fundamental functional spaces that take into account the incompressibility condition (4)2 (cf. [87]):
@@
The operator $A:D(A)\subset H\to H$ is defined as $Au=-P_{L}\Delta{\bf{u}}$, where $P_{L}$ is the Helmholtz–Leray projection from $L^{2}(\mathbb{T}^{3};\mathbb{R}^{3})$ onto $H$ and $D(A)=H^{2}(\mathbb{T}^{3};\mathbb{R}^{3})\cap V$. It is well-known that $A^{-1}$ is well-defined from $H$ to $D(A)$. It is also self-adjoint. By the spectral theory of compact self-adjoint operators in a Hilbert space, there exists a sequence of eigenvalues $0<\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{m}\leq\cdots$, $\lambda_{m}\to\infty$ as $m\to\infty$ and an orthonormal basis $\{e_{i}\}_{i=1}^{\infty}$ such that $Ae_{i}=\lambda_{i}e_{i}$. The fractional power $A^{\alpha}$ of $A$, $\alpha\geq 0$, is simply defined by
@@
We also note that $D(A^{\alpha})\subset H^{2\alpha}(\mathbb{T}^{3};\mathbb{R}^{3})$. For the definition of $H^{2\alpha}(\mathbb{T}^{3};\mathbb{R}^{3})$, we refer to [78, Definition 1.10].
2.2 Stochastic framework
@@
$\mathfrak{B}:=\left(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P},\left(W(t)\right)_{t\geq 0}\right),$ (5)
@@
$\left(W(t)\right)_{t\geq 0}$ takes value in $D(A^{\delta})$ for some $\delta>0$ with a covariance operator $\mathcal{L}$. (6)
Here $\mathcal{L}\in\mathcal{B}(D(A^{\delta}))$ and is non-negative and of trace class. Let $\{b_{i}\}_{i\geq 1}$ be the orthonormal basis of $D(A^{\delta})$ consisting of eigenvectors of $\mathcal{L}$ subject to non-negative eigenvalues $\{\gamma_{i}\}_{i\geq 1}$. Then we have the following expansion (cf. [74, Theorem 4.20])
@@
$W(t)\in C^{\beta}\left([0,T];D(A^{\delta})\right)\quad\mathbb{P}-a.s..$ (7)
For the purpose of obtaining a key preliminary estimation later on (see Lemma 1 in Section 3.2), we will need to further assume that $\delta\geq\frac{3}{4}$. So this assumption together with (6) yields the assumption that we require on the noise:
$\left(W(t)\right)_{t\geq 0}$ takes value in $D(A^{\delta})$ for some $\delta>3/4$ with a covariance operator $\mathcal{L}$. (8)
3 Deterministic and martingale suitable weak solutions
In order to define suitable weak solutions that are deterministic and are martingale, we follow the approach of [26, 79], and split the SEL into the sum of a linear stochastic Stokes equation and a nonlinear modified Ericksen-Leslie equation.
3.1 A split-up scheme of the SEL
@@
We will assume throughout the paper the following assumptions on the initial data
$({{\bf{u}}}_{0},{{\bf{d}}}_{0})\in H\times H^{1}(\mathbb{T}^{3};\mathbb{R}^{3}),\quad|{{\bf{d}}}_{0}(x)|=1\text{ for all }x\in\mathbb{T}^{3}.$ (9)
@@
As a preparation to introduce notions of deterministic and stochastic suitable weak martingale solutions, we will first split the solutions ${{\bf{u}}}$ and $P$ properly. We introduce new variables ${{\bf{u}}}={{\bf{z}}}+{{\bf{v}}}$, $P={Q}+{\pi}$, where $({{\bf{z}}},{Q})$ solves the linear stochastic Stokes system as follows
@@
$\left\{\begin{array}[]{l}{\,\mathrm{d}}{{\bf{z}}}-\Delta{{\bf{z}}}{\,\mathrm{d}}t+\nabla{Q}{\,\mathrm{d}}t={\,\mathrm{d}}W,\\ \mathop{\rm div}\nolimits{{\bf{z}}}=0,\\ {{\bf{z}}}(0)=0.\end{array}\right$. (10)
The triple $({{\bf{v}}},{\pi},{{\bf{d}}})$ solves the modified Ericksen–Leslie equations as follows
@@
$\begin{cases}&\partial_{t}{{\bf{v}}}-\Delta{{\bf{v}}}+({{\bf{z}}}+{{\bf{v}}})\cdot\nabla({{\bf{z}}}+{{\bf{v}}})+\nabla{\pi}\\ &\quad\quad\quad\quad\quad=-\mathop{\rm div}\nolimits\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}-F({{\bf{d}}}){\rm I}_{3}\right),\\ &\mathop{\rm div}\nolimits{{\bf{v}}}=0,\\ &\partial_{t}{{\bf{d}}}-\Delta{{\bf{d}}}+({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{d}}}=-f({{\bf{d}}}),\\ &{{\bf{v}}}(0)={{\bf{u}}}_{0},{{\bf{d}}}(0)={{\bf{d}}}_{0}.\end{cases}$ (11)
Remark 1.
@@
Thanks to the incompressibility condition (11)2, the right-hand side of (11)1 is is equivalent to the deterministic term on the right-hand side of (4)1. Specifically, for the extra term, we observe:
$\mathop{\rm div}\nolimits\left(\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}+F({{\bf{d}}}){\rm I}_{3}\right)=\nabla\left(\frac{1}{2}|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right),$
@@
which can be treated as part of the pressure gradient. Moreover, this observation, together with direct computations, yields
@@
$\mathop{\rm div}\nolimits\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}-F({{\bf{d}}}){\rm I}_{3}\right)=-\nabla{\bf{d}}\cdot(\Delta{\bf{d}}-f({\bf{d}})).$ (12)
As a result, for computational convenience, we shall use the right-hand-side of the above equation to replace the right-hand side of of (11)1.
3.2 Definitions of suitable weak solutions
In order to define what it means for the modified Ericksen-Leslie equation (11) to have a weak solution, we need a preliminary estimation of the solution to the linear stochastic Stokes system (10). We first derive a result that ensures the boundedness of ${{\bf{z}}}(t)$, the stochastic process that solves (10). To achieve this goal, we need to utilize the following result, which is also in [26] (see [26, Lemma 2.2]).
Lemma 1 (An almost-sure bound of the linear stochastic Stokes system).
With the assumption on the noise as in (8), we have
${{\bf{z}}}(\omega)\in L_{\rm loc}^{\infty}(\mathbb{T}^{3}\times[0,\infty))\mbox{ for $\displaystyle\mathbb{P}$- a.s. $\displaystyle\omega$}.$ (13)
The proof of Lemma 1 is provided in Appendix A.
With Lemma 1 in hand, now we are ready to introduce the notions of deterministic and martingale suitable weak solutions to SEL.
Definition 1.